OpenStudy (anonymous):

You are familiar with the following types of factoring: factoring out the Greatest Common Factor (GCF) factoring by grouping factoring trinomials of the form x2 + bx + c and ax2 + bx + c As you know, you need to know the first two types of factoring listed above in order to be successful in factoring trinomials of the form ax2 + bx + c. Part 1: In your own words, explain how a trinomial of the form 2x2 + 13x + 15 can be turned into a four term polynomial suitable for factoring by grouping. Use complete sentences.

OpenStudy (anonymous):

Part 2: If you were an Algebra 1 instructor and were creating a test on factoring trinomials of the form ax2 + bx + c, what do you think would be the easiest way to create a trinomial that can be factored? Provide one unique example.

OpenStudy (anonymous):

for part 1 , do you have any ideas?

OpenStudy (anonymous):

No I'm not so good at these... :/

OpenStudy (anonymous):

Alright so in the form \[ax^2+bx+c=0\] if you if you multiply a by c which in this case you get \[2*15=30\] now we need to find the factors of this number that add up to b or 13 this would be \[10*3=30\] \[10+3=13\] now you just split the bx into the factors 10 and a 3 \[2x^2+10x+3x+15=0\]

OpenStudy (anonymous):

group the first two terms, and the last 2 \[(2x^2+10x)+(3x+15)=0\]

OpenStudy (anonymous):

take the GCF of each parenthesis \[2x(x+5)+3(x+5)=0\] if you get the same thing in parenthesis, your factored polynomial whats inside the parenthesis and whats outside so \[(2x+3)(x+5)=2x^2+13x+15\]

OpenStudy (anonymous):

Thank you so much!!! :)

OpenStudy (anonymous):

make sense? so part 2 what do you think you should do?